On January 31, a snowstorm damaged the office of a small business, and some of the accounting information stored in the computer’s memory was lost. The following information pertaining to January activities was retrieved from other sources:

Direct materials purchased $18,000

Work in process—beginning inventory 2,000

Direct materials—beginning inventory 6,000

Direct materials—ending inventory 10,000

Finished goods—beginning inventory 12,000

Finished goods—ending inventory 2,500

Sales 60,000

Manufacturing overhead and direct labour incurred 22,000

Gross profit percentage based on net sales 40 %

a.What was the cost of direct materials used in January?

b.What amount of work in process inventory was transferred out to finished goods during January?

c. Assume that $20,000 of direct materials was used in January and that the cost of goods available for sale in January amounted to $40,000. What did the ending work in process inventory amount to?
Answer #1

a) Direct material used = Beginning inventory+Purchase-Ending inventory = 6000+18000-10000 = 14000

b) Work in process transferred out to finished goods =

Cost of goods sold = 60000*60% = 36000

Work in process transferred out to finished goods = 36000+2500-12000 = 26500

c) Cost of goods manufactured = 40000-12000 = 28000

Ending work in process = 2000+20000+22000-28000 = 16000